# Circle packing and Riemann uniformization of random triangulations in an ergodic scale-free environment

### 06 Mar 2026 v1

@@ 3.2 Convergence of random walk with Dubejko weights on cell configurations
We begin with the convergence of random walks on cell configurations in [GMS22]. Let $\beta_{1}:[0,T_{\beta_{1}}]\to\mathbb{C}$ and $\beta_{2}:[0,T_{\beta_{2}}]\to\mathbb{C}$ be continuous curves defined on possibly different time intervals. We define
Theorem 3.5 (Theorem 3.10 of [GMS22]).
Corollary 3.6.
[…]
Proof.
@@ 3.3.1 Vertex extremal length and macroscopic circles
Recall the notion of vertex extremal length introduced in Section 3.3.1. Let $\widehat{\mathcal{H}}$ be a planar embedding of the associated map of $\mathcal{H}$ that is close to $\mathcal{H}$ in the sense of Definition 1.5, with vertex set $\{z_{H}:H\in\mathcal{H}\}$. For each $H\in\mathcal{H}$, let $\Gamma_{H}$ be the collection of finite paths on $\widehat{\mathcal{H}}$ starting from $z_{H}$ and making a simple loop around $z_{H_{0}}$.
@@
Without loss of generality assume $q_{0}>300$ and set $q=q_{0}/72$. We will adapt the proof of [GJN19, Theorem 1.4] together with the input from Section $2.$ For $c>0$ and any square $S$, define the event $E_{1}(c,S)$ where
@@
and therefore $\mathrm{area}({m^{*}})\leq 16c_{0}(n-1)$. On the other hand, using the definition of $E_{k,2}(\varepsilon)$, $H^{*}\subseteq S_{1}^{*}$ and $H_{0}\cap S_{n}^{*}=\emptyset$. For each path $\ell\in\Gamma_{H^{*}}$, using $H\cap H^{\prime}\neq\emptyset$ for $H\sim H^{\prime}$ for all large enough $k$, $\cup_{H\in\ell}H$ must cross all the regions $S_{j}^{*}\backslash S_{j-1}^{*}$, which implies
@@
Our proof of Theorem 1.7 is mostly based on the convergence of the same random walk on planar maps with different embeddings to the Brownian motion as in Proposition 3.8 and Proposition 3.9, as well as the consideration of wrapping around events for 2D Brownian motion. In the rest of this section, we use the following notation. For a continuous-time walk $(X_{t})_{t\geq 0}$ on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ and a bounded set $A\subseteq\mathbb{C}$ where $z\notin A$, we write $\mathbb{P}^{z}$ for the law of $(X_{t})_{t\geq 0}$ started from $z$, and $\tau_{A}$ for the first time $T$ when the trace of $(X_{t})_{0\leq t\leq T}$ separates AA from infinity. We write $\tau_{w}$ instead of $\tau_{\{w\}}$ if $A=\{w\}$.
@@
Let $\widehat{X}^{z,\rm circle,\widetilde{\varepsilon}}$ be the linearly interpolated random walk on $\widetilde{\varepsilon}\mathcal{P}$ described in Proposition 3.9, and let $\widehat{X}^{z,\varepsilon}$ be interpolated random walk on $\varepsilon\mathbf{A}\mathcal{H}$ as in Theorem 3.5 using the points $\{\mathbf{A}z_{H}:H\in\mathcal{H}\}$, except that we interpolate along the edges $\{\mathbf{A}\gamma_{H,H^{\prime}}:H\sim H^{\prime}\}$. Note that $f^{\varepsilon}(\widehat{X}^{z,\varepsilon})$ and $\widehat{X}^{z,\rm circle,\widetilde{\varepsilon}}$ can be coupled to have the same trace. Indeed, this coupling exists since $\varepsilon\mathcal{H}$ and $\widetilde{\varepsilon}\mathcal{P}$ can be viewed as two embeddings of the associated map for $\mathcal{H}$ and have identical conductance, and we work on this coupling. Then by Proposition 3.8 and Proposition 3.9, as $\varepsilon\to 0$ the conditional law of $\widehat{X}^{z,\varepsilon}$ and $\widehat{X}^{z,\rm circle,\widetilde{\varepsilon}}$ given $\varepsilon\mathcal{H}$ and $\widetilde{\varepsilon}\mathcal{P}$ converges to a constant multiple of planar Brownian motion with respect to the topology on curves induced in (3.7). Throughout the proof, we write $\mathbb{P}_{\rm BM}$ for the law of standard Brownian motion, $\mathbb{P}_{\varepsilon}$ for the conditional law of $\widehat{X}^{\varepsilon}$ given $\varepsilon\mathcal{H}$, and $\widetilde{\mathbb{P}}_{\widetilde{\varepsilon}}$ for the conditional law of $\widehat{X}^{\rm circle,\widetilde{\varepsilon}}$ given $\widetilde{\varepsilon}\mathcal{P}$. Also let $K$ be a fixed compact subset of $\mathbb{C}$ containing 0 and 1, and $d_{K}=\mathrm{diam}(K)$. We will verify that $\{f^{\varepsilon}\}$ is uniformly bounded and equicontinuous on compacts.

### 06 May 2026 v2

@@ 2.2 Ergodic averages of cell configurations using dyadic systems
For a square $S$, we define a disk triangulation $\mathcal{M}(S)$ using cells in $\mathcal{H}(S)$ as follows. Let $\mathcal{M}$ be the associated map of the cell configuration $\mathcal{H}$. For every $a>0$, let $\mathring{S}^{a}$ be the square concentric to $S$ with side length $a$. We work on the event that $\mathcal{H}(\mathring{S}^{a})$ is connected; otherwise we set $\mathcal{M}(S;a)=\emptyset$. Let $\mathcal{M}_{0}(S;a)\subseteq\mathcal{M}$ be the submap formed by the union of faces of $\mathcal{M}$ with at least one of the vertices in $\mathcal{H}(\mathring{S}^{a})$. Note that when viewed as a graph, $\mathcal{M}_{0}(S;a)$ is 2-connected since $\mathcal{H}(\mathring{S}^{a})$ is connected. In fact, one can verify that, for any infinite plane triangulation $\mathcal{T}$ and a connected subgraph $\mathcal{G}_{1}\subseteq\mathcal{G}$, if we let $\mathcal{G}_{2}$ be the subgraph $\mathcal{G}$ formed by all the vertices in $\mathcal{G}_{1}$ and their neighbors, as well as all the edges connecting these vertices, then $\mathcal{G}_{2}$ is 2-connected. We further let $\mathcal{M}(S;a)$ be the smallest simply connected submap of $\mathcal{M}$ containing $\mathcal{M}_{0}(S;a)$ as a submap. Then $\mathcal{M}(S;a)$ is a disk triangulation. Let $\mathcal{M}(S)$ be the largest submap of the form $\mathcal{M}(S;a)$ for some $a\in(0,|S|)$ such that for each vertex in $\mathcal{M}(S;a)$, its cell is contained in the interior of $S$. Note that these objects are well defined since $\mathcal{H}$ is locally finite. We further write $M(S)$ for the Riemann surface associated with $\mathcal{M}(S)$. See Figure 3 for an illustration.
Lemma 2.8.
Almost surely, for each given $\varepsilon>0$ and $a\in(0,1)$, for large enough $k$, for each square $S\subseteq[-2^{k},2^{k}]^{2}$ of side length at least $a2^{k}$, $\mathcal{M}(S;(1-\varepsilon)|S|)\subseteq\mathcal{M}(S)$.
Proof.
[…]
Let $\mathcal{H}$ be the cell configuration in the setting of Theorem 1.6 or Theorem 1.7. We introduce a function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$ below that will later be used in Section 2.3 and Section 4. The function $\phi_{0}$ sends each vertex on $M(\mathcal{H})$ to a point in the associated cell of $\mathcal{H}$ and is otherwise piecewise linear. More precisely, the function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$ is defined as follows. Given a realization of $\mathcal{H}$, for each cell $H\in\mathcal{H}$, we independently sample a point $\widetilde{c}(H)$ according to the probability measure proportional to the Lebsegue measure on $H$. For each triangular face $\Delta$ in $M(\mathcal{H})$ formed by the cells $H_{1},H_{2},H_{3}\in\mathcal{H}$, consider the triangle $\Delta^{\prime}$ in $\mathbb{C}$ formed by connecting $\widetilde{c}(H_{1}),\widetilde{c}(H_{2}),\widetilde{c}(H_{3})$ using straight line segments, and define $\phi_{0}|_{\Delta}:\Delta\to\Delta^{\prime}$ to be the unique linear function sending the vertices in $\Delta$ to their counterpart in $\Delta^{\prime}$. This defines a continuous and piecewise linear function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$. See Figure 4 for an illustration.
We end this subsection with the following topological properties on $\phi_{0}$.
Lemma 2.9.
We have $\phi_{0}(M(\mathcal{H}))=\mathbb{C}$.
Proof.
Pick $k$ large enough such that by Definition 1.4, the 4 subgraphs of $\mathcal{H}$ formed by the cells intersecting the 4 sides of $S_{k}$ are connected. Since a cell intersecting a vertex of $S_{k}$ is contained in (at least) two of these subgraphs, we get that the set of cells intersecting a side of $S_{k}$ forms a connected graph. By further applying Lemma 2.6, we can find a chain of cells $H_{1}\sim H_{2}\sim\cdots\sim H_{n}\sim H_{1}$ in this connected graph such that if we connect $\widetilde{c}(H_{j})$ and $\widetilde{c}(H_{j+1})$ using straight line segments for j=1,,nj=1,...,n with $n+1$ identified with 1, then this creates a loop $\gamma_{n}$ surrounding the square $\widetilde{S}_{k}$ concentric to $S_{k}$ with side length $|S_{k}|/2$, and the winding number of $\gamma_{n}$ around each point in $\widetilde{S}_{k}$ is equal to 1. If we consider the vertices for $H_{1},...,H_{n}$ on the surface $M(\mathcal{H})$, then one can find a loop $\widetilde{\gamma}_{n}$ on $M(\mathcal{H})$ connecting these vertices with straight line segments such that $\phi_{0}(\widetilde{\gamma}_{n})=\gamma_{n}$. Since $M(\mathcal{H})$ is simply connected, we can continuously shrink $\widetilde{\gamma}_{n}$ to a single point $x$. If there exists a point $w\in\widetilde{S}_{k}\backslash\phi_{0}(M(\mathcal{H}))$, then under $\phi_{0}$ image, all of these shrinking loops will have winding number 1 around $w$. On the other hand, for loops sufficiently close to $x$, its image under $\phi_{0}$ has winding number 0 around $w$ since $w\notin\phi_{0}(M(\mathcal{H}))$. This is a contradiction, which implies that almost surely for large enough $k$, $\widetilde{S}_{k}\subseteq\phi_{0}(M(\mathcal{H}))$, which completes the proof. ∎
Lemma 2.10.
For every compact set $K\subseteq\mathbb{C}$, $\phi_{0}^{-1}(K)$ has compact closure in $M(\mathcal{H})$.
Proof.
By Lemma 2.6, a.s. for all large enough $k$, any cells in $\mathcal{H}([-2^{k},2^{k}]^{2})$ has diameter less than $2^{k-10}$, and the convex hull of the union of any three cells $H_{1}\sim H_{2}\sim H_{3}$ in $\mathcal{H}([-2^{k},2^{k}]^{2}\backslash[-2^{k-2},2^{k-2}]^{2})$ does not intersect $K$. This gives the claim. ∎
Lemma 2.11.
Almost surely for large enough $k$, the following is true. For any cells $H_{1},...,H_{n}\in\mathcal{H}(\partial[-2^{k},2^{k}]^{2})$ such that $H_{1}\sim H_{2}\sim...\sim H_{n}\sim H_{1}$ form a simple loop $\gamma_{n}$ on $M(\mathcal{H})$ by drawing a straight line between the vertices representing adjacent cells, if $\phi_{0}(\gamma_{n})$ has winding number 1 around 0, then the loop $\gamma_{n}$ encloses $M([-2^{k},2^{k}]^{2})$.
Proof.
From the definition, $M([-2^{k},2^{k}]^{2})$ has disk topology and is disjoint from $\gamma_{n}$. Suppose $k$ is large enough such that by Lemma 2.10, $\phi_{0}^{-1}(0)\subseteq M([-2^{k-1},2^{k-1}]^{2})$. If $\gamma_{n}$ does not enclose $M([-2^{k},2^{k}]^{2})$, then one can continuously shrink $\gamma_{n}$ to a single point $x$ on $M(\mathcal{H})$ such that all the loops are disjoint from $M([-2^{k},2^{k}]^{2})$. In particular, under the image of $\phi_{0}$, all of these shrinking loops will have winding number 1 around 0. On the other hand, $x\notin M([-2^{k},2^{k}]^{2})$ and $\phi_{0}(x)\neq 0$. This is a contradiction. ∎
2.3 Convergence of random walks on cell configurations
In order to prove Theorem 1.6 and Theorem 1.7, we will need that for each large square $S$, $\mathcal{H}(S)$ is close to some planar embedding of the map $\mathcal{M}(S)$. To achieve this, in Proposition 2.15 below, we will use the the convergence of random walks on cell configurations in [GMS22] to prove that for large squares $S$, $\mathcal{H}(S)$ is close to the Tutte embedding of $\mathcal{M}(S)$.
Let $\beta_{1}:[0,T_{\beta_{1}}]\to\mathbb{C}$ and $\beta_{2}:[0,T_{\beta_{2}}]\to\mathbb{C}$ be continuous curves defined on possibly different time intervals. We define
Theorem 2.12 (Theorem 3.10 of [GMS22]).
Corollary 2.13.
[…]
Proof.
As studied in [GMS21], one main application of Theorem 2.12 is the convergence of random planar maps under the Tutte embedding. Let $\mathcal{M}_{0}$ be a disk triangulation and $v\in\mathcal{V}\mathcal{M}_{0}$ be an interior vertex. The Tutte embedding of $\mathcal{M}_{0}$ in $\mathbb{D}$ centered at $v$ is defined as follows. Enumerate the boundary vertices of $\mathcal{M}_{0}$ in counterclockwise order as $\{y_{1},...,y_{n}\}$ with $y_{1}$ chosen arbitrarily. For j=1,,nj=1,...,n we let $\mathfrak{p}(y_{j})$ be the probability that a simple random walk started from $v$ first hits the boundary at a vertex in the boundary arc $\{y_{1},...,y_{j}\}$. The boundary vertices $y_{1},...,y_{n}$ are then mapped in counterclockwise order to the unit circle $\partial\mathbb{D}$ such that $\psi(y_{j}):=e^{2\pi i\mathfrak{p}(y_{j})}$. Then the hitting probability of the random walk started from $v$ approximates the uniform measure on the unit circle. We map the interior vertices of $\mathcal{M}_{0}$ into the unit disk via the discrete harmonic extension of the boundary values, i.e., for any interior vertex $v$,
$\psi(v)=\frac{1}{\#\{w\in\mathcal{V}\mathcal{M}_{0}:w\sim v\}}\sum_{w\sim v}\psi(w).$ (2.11)
Let $\psi:\mathcal{V}\mathcal{M}_{0}\to\mathbb{D}$ be the discrete harmonic function with boundary data as prescribed above. For each edge $e$ of $\mathcal{M}_{0}$ connecting $v_{1}$ and $v_{2}$, we also define $\psi(e)$ to be the straight line segment joining $\psi(v_{1})$ and $\psi(v_{2})$. This defines the Tutte embedding [^1] Since $\mathcal{M}_{0}$ potentially has multiple edges, the Tutte embedding might not be a planar embedding as defined at the beginning of Section 1.1 since $\psi$ might be non-injective. On the other hand, the embedding in (2.12) below will be an actual planar embedding. of the map $\mathcal{M}_{0}$ within the unit disk $\overline{\mathbb{D}}$. Note that under this drawing, if $e_{1},...,e_{k}$ are edges with the same endpoints (i.e., these edges form a multiple edge), then $\psi(e_{1})=...=\psi(e_{k})$, and vertices enclosed by the edges $e_{1},...,e_{k}$ on the map $\mathcal{M}_{0}$ are mapped to points on the line segment $\psi(e_{1})$. Furthermore, the drawing $\psi$ is planar in the following sense. Let $e_{1}$ and $e_{2}$ be edges that are not contained within multiple edges of $\mathcal{M}_{0}$, i.e., they are not part of a multiple edge or are enclosed by one. If $e_{1}$ and $e_{2}$ have different endpoints, $\psi(e_{1})\cap\psi(e_{2})=\emptyset$, and if $e_{1}$ and $e_{2}$ has one common endpoint $v$, then $\psi(e_{1})\cap\psi(e_{2})=\psi(v)$.
Next we prove that for large $k$, the embedding of $\mathcal{M}([-2^{k},2^{k}]^{2})$ via $\mathcal{H}$ is close to some planar embedding of $\mathcal{M}([-2^{k},2^{k}]^{2})$. We will typically use the Tutte embedding here as the planar embedding. However, the embedded boundary of $\mathcal{M}([-2^{k},2^{k}]^{2})$ via $\mathcal{H}$ only has close Hausdorff distance, not curve distance (2.8), to $\partial[-2^{k},2^{k}]^{2}$, which is not sufficient. Below we will construct a map $\mathcal{M}_{k,\ell}$ which contains and is slightly larger than $\mathcal{M}([-2^{k},2^{k}]^{2})$ and work on that map instead. Recall the function $\phi_{0}$ defined after Lemma 2.8.
Lemma 2.14.
Let $\ell>0$. Almost surely, for all large enough $k$, there exists a submap $\mathcal{M}_{k,\ell}\subseteq\mathcal{M}$ with the following properties:
  1. (i)
    $\mathcal{M}_{k,\ell}$ has disk topology, and for each vertex $v$ on its boundary, its cell in $\mathcal{H}$ intersects $\partial[-2^{k},2^{k}]^{2}$;
  2. (ii)
    $\mathcal{M}_{k,\ell}$ contains $\mathcal{M}([-2^{k},2^{k}]^{2})$ as a submap;
  3. (iii)
    Let $M_{k,\ell}\subseteq M(\mathcal{H})$ be the Riemann surface generated from $\mathcal{M}_{k,\ell}$. Then the curve distance (2.8) between $\phi_{0}(\partial M_{k,\ell})$ and $\partial[-2^{k},2^{k}]^{2}$ is less than $2^{k-\ell}$.
Proof.
Following Definition 1.4 and Lemma 2.6, let $k$ be large enough such that for any line segment $L\subseteq[-2^{k+1},2^{k+1}]$ of length at least $2^{k-\ell-4}$, $\mathcal{H}(L)$ is connected, and the maximal diameter of cells in $\mathcal{H}([-2^{k+1},2^{k+1}]^{2})$ is less than $2^{k-2\ell-10}$. We partition $\partial[-2^{k},2^{k}]^{2}$ into line segments $L_{1},...,L_{2^{\ell+2}}$ of length $2^{k-\ell}$, such that $L_{1}=[-2^{k},-2^{k}+2^{k-\ell}]\times\{-2^{k}\}$, and $L_{1},...,L_{2^{\ell+2}}$ are aligned counterclockwise. For $j=1,...,2^{\ell+2}$ with $2^{\ell+2}+1$ identified with 1, let $\mathring{L_{j}}$ be the line segments concentric to $L_{j}$ with length $|L_{j}|/2$, and let $L_{j,j+1}$ be the union of $\mathring{L_{j}},\mathring{L}_{j+1}$ and the middle component of $(L_{j}\cup L_{j+1})\backslash(\mathring{L_{j}}\cup\mathring{L}_{j+1})$. We pick a cell $H_{j}\in\mathcal{H}(\mathring{L_{j}})$. By the line connectivity property above, we can find a simple path $\widetilde{\gamma}_{k,j}$ on $\mathcal{H}(L_{j,j+1})$ connecting $H_{j}$ and $H_{j+1}$. The intersection between $\widetilde{\gamma}_{k,j}$ and $\widetilde{\gamma}_{k,j+1}$ are in $\mathcal{H}(\mathring{L}_{j+1})$, and therefore we can erase all the loops in $\cup_{j}\widetilde{\gamma}_{k,j}$ that are made by vertices in $\mathcal{H}(\mathring{L}_{j+1})$ for each $j$ to obtain a simple loop $\gamma_{k}$. Let $\mathcal{M}_{k,\ell}\subseteq\mathcal{M}$ be the planar map formed by $\gamma_{k}$ and the vertices enclosed by $\gamma_{k}$. Then properties (i) and (iii) are immediate from construction, while property (ii) is a consequence of Lemma 2.11. ∎
We are now ready to prove that the embedding of $\mathcal{M}([-2^{k},2^{k}]^{2})$ via $\mathcal{H}$ is close to planar.
Proposition 2.15.
Let $\mathcal{H}$ be a cell configuration in the context of Theorem 1.6 or Theorem 1.7. There exists a collection of points $\{z_{H}^{k}:H\in\mathcal{V}\mathcal{M}([-2^{k},2^{k}]^{2})\}$ and simple curves $\{\gamma_{e}^{k}:e\in\mathcal{E}\mathcal{M}([-2^{k},2^{k}]^{2})\}$ forming a planar embedding of $\mathcal{M}([-2^{k},2^{k}])$, such that almost surely,
$\frac{1}{2^{k}}\max_{H\in\mathcal{V}\mathcal{M}([-2^{k},2^{k}]^{2})}\max_{e}\operatorname{dist}(\gamma_{e},H)\to 0,$ (2.12)
Proof.
By setting the conductance to be equal to 1 for all edges, the moment bound (2.6) implies the moment bound (2.10), so the conclusion in Theorem 2.12 holds. Without loss of generality assume that the matrix $\Sigma$ there is the identity matrix; otherwise we choose a deterministic matrix $\mathbf{A}$ such that the random walk on $\mathbf{A}\mathcal{H}$ converges to the standard planar Brownian motion in the sense of Theorem 2.12.
Let $k,\ell>0$, the map $\mathcal{M}_{k,\ell}$ be as in Lemma 2.14, and $L_{1},...,L_{2^{\ell+2}}$ be as in the proof of Lemma 2.14. Let $X^{z}$ and $\widehat{X}^{z}$ be the random walks in Theorem 2.12 started from a cell in the interior of $\mathcal{M}_{k,\ell}$, and let $\widehat{H}$ be the cell where the walk $X^{z}$ first hits $\partial\mathcal{M}_{k,\ell}$. From our construction, if $\widehat{H}$ is on $\mathcal{H}(L_{1}\cup...\cup L_{j})$, then the walk $\widehat{X}^{z}$ hits the $2^{k-\ell+1}$-neighborhood of $L_{1}\cup...\cup L_{j}$ before hitting $L_{j+2}\cup...\cup L_{2^{\ell+2}-1}$. Likewise, if $\widehat{H}$ is on $\mathcal{H}(L_{j+1}\cup...\cup L_{2^{\ell+2}})$, then the walk $\widehat{X}^{z}$ hits the $2^{k-\ell+1}$-neighborhood of $L_{j+1}\cup...\cup L_{2^{\ell+2}}$ before hitting $L_{1}\cup...\cup L_{j-1}$. Further applying Theorem 2.12, a.s. we send $\ell\to\infty$ sufficiently slower than $k\to\infty$, for each $a\in(0,1)$ and $z\in[-a2^{k},a2^{k}]^{2}$, the conditional probability given $\mathcal{H}$ of the walk $X^{z}$ leaves $\mathcal{M}_{k,\ell}$ at a vertex on $\mathcal{H}(L_{1}\cup...\cup L_{j})$ is close to the harmonic measure on $\partial[-2^{k},2^{k}]^{2}$ as seen from $z$ of $L_{1}\cup...\cup L_{j}$. Therefore when embedded into $\mathbb{D}$, by the conformal invariance of Brownian motion, if we take a conformal map $\varphi_{0}:[-1,1]^{2}\to\mathbb{D}$ fixing 0, then up to a rotation, $2^{-k}\varphi_{0}(H)$ is close to the vertex for $H$ under the Tutte embedding of $\mathcal{M}_{k,\ell}$ centered at $H_{0}$. We can further consider the image of the Tutte embedding for $\mathcal{M}_{k,\ell}$ under $2^{k}\varphi_{0}^{-1}$ to generate the points and curves $\{\widetilde{z}_{H}^{k}\}$ and curves $\{\widetilde{\gamma}_{e}\}$ where (2.12) holds, and further separate the curves $\widetilde{\gamma}_{e}$ a little for multiple edges and embed the components contained within these multiple edges in an arbitrary way to get an actual planar embedding where (2.12) holds.
@@ 3.2 Convergence of random walk with Dubejko weights on cell configurations
We start with the following immediate consequence of Theorem 2.12 and Theorem 2.1.
@@ 3.3.1 Vertex extremal length and macroscopic circles
Recall the notion of vertex extremal length introduced in Section 3.3.1. For each $H\in\mathcal{H}$, let $\Gamma_{H}$ be the collection of finite paths on $\mathcal{M}$ starting from $H$ and making a simple loop around $H_{0}$}.
@@
Without loss of generality assume $q_{0}>300$ and set $q=q_{0}/72$. We will adapt the proof of [GJN19, Theorem 1.4] together with the input from Section 2.2. For $c>0$ and any square $S$, define the event $E_{1}(c,S)$ where
@@
and therefore $\mathrm{area}({m^{*}})\leq 16c_{0}(n-1)$. On the other hand, using the definition of $E_{k,2}(\varepsilon)$, $H^{*}\subseteq S_{1}^{*}$ and $H_{0}\cap S_{n}^{*}=\emptyset$. For each path $\ell\in\Gamma_{H^{*}}$, using $H\cap H^{\prime}\neq\emptyset$ for $H\sim H^{\prime}$ for all large enough $k$, together with the almost planarity condition for $\mathcal{M}([-2^{k+3},2^{k+3}]^{2})$ from Proposition 2.15, $\cup_{H\in\ell}H$ must cross all the regions $S_{j}^{*}\backslash S_{j-1}^{*}$, which implies
@@
Our proof of Theorem 1.6 is mostly based on the convergence of the same random walk on planar maps with different embeddings to the Brownian motion as in Proposition 3.6 and Proposition 3.7, as well as the consideration of wrapping around events for 2D Brownian motion. In the rest of this section, we use the following notation. For a continuous-time walk $(X_{t})_{t\geq 0}$ on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ and a bounded set $A\subseteq\mathbb{C}$ where $z\notin A$, we write $\mathbb{P}^{z}$ for the law of $(X_{t})_{t\geq 0}$ started from $z$, and $\tau_{A}$ for the first time $T$ when the trace of $(X_{t})_{0\leq t\leq T}$ separates $A$ from infinity. We write $\tau_{w}$ instead of $\tau_{\{w\}}$ if $A=\{w\}$. For $\ell>10$, we write $\sigma_{\ell}$ for the first time when the walk exits $2^{\ell-1},2^{\ell-1}]^{2}$.
@@
Let $\widehat{X}^{z,\mathrm{circle},k,\ell}$ be the linearly interpolated random walk on $\widetilde{\varepsilon}_{k}\mathcal{P}$ described in Proposition 3.7, and let $\widehat{X}^{z,k,\ell}$ be interpolated random walk on $2^{-k}\mathbf{A}\mathcal{H}$ as in Theorem 2.12, except that for $H\in\mathcal{V}\mathcal{M}([-2^{k+\ell},2^{k+\ell}]^{2})$ we use the points $\{\mathbf{A}z_{H}^{k+\ell}\}$ and for $e\in\mathcal{E}\mathcal{M}([-2^{k+\ell},2^{k+\ell}]^{2})$ we interpolate along the edges $\{\mathbf{A}\gamma_{e}^{k+\ell}:e\in\mathcal{E}\mathcal{M}([-2^{k+\ell},2^{k+\ell}]^{2})\}$.Note that by Lemma 2.8, almost surely for large enough $k$, the traces of $f^{k}(\widehat{X}^{z,k,\ell})$ and $\widehat{X}^{z,\mathrm{circle},k,\ell}$ within $f^{k,\ell}([-2^{\ell-1},2^{\ell-1}]^{2})$ can be coupled to be the same. Indeed, this coupling exists since $\varepsilon\mathcal{H}$ and $\widetilde{\varepsilon}\mathcal{P}$ can be viewed as two embeddings of the associated map for $\mathcal{H}$ and have identical conductance, and we work on this coupling. Then by Proposition 3.6 and Proposition 3.7, as $k\to\infty$ the conditional law of $\widehat{X}^{z,k,\ell}$ and $\widehat{X}^{z,\mathrm{circle},k,\ell}$ given $\varepsilon_{k}\mathcal{H}$ and $\widetilde{\varepsilon}_{k}\mathcal{P}$ converges to a constant multiple of planar Brownian motion with respect to the topology on curves induced in (2.9). Throughout the proof, we write $\mathbb{P}_{\rm BM}$ for the law of standard Brownian motion, $\mathbb{P}_{k,\ell}$ for the conditional law of $\widehat{X}^{k,\ell}$ given $\varepsilon\mathcal{H}$, and $\widetilde{\mathbb{P}}_{k,\ell}$ for the conditional law of $\widehat{X}^{\mathrm{circle},k,\ell}$ given $\widetilde{\varepsilon}_{k}\mathcal{P}$. Fix $\widetilde{\ell}>10$, let $K=[-2^{\widetilde{\ell}-5},2^{\widetilde{\ell}-5}]$ and $d_{K}=2^{\widetilde{\ell}-4}$. We will verify that $\{f^{k,\ell}{\,:\,k\in\mathbb{N},\ell>\widetilde{\ell}+1}\}$ is uniformly bounded and equicontinuous on $K$ for any fixed $\ell$, and we will later send $\widetilde{\ell}\to\infty$.